# Redmond, WA %TODO␣FINAL:␣Replace␣with␣your␣institution␣list.www.retrocausal.ai

### v5

@@ 1 Introduction
00footnotetext: indicates joint first author.
{huy,ahmed,murad,hassan,andrey,zeeshan}@retrocausal.ai.
Figure 1: (a) The state-of-the-art method in video-based self-supervised learning, i.e., CASA, uses 3D skeleton-based temporal video alignment as a pretext task and performs 3D skeleton augmentations. (b) Our approach relies on 2D skeleton-based temporal video alignment and conducts 2D skeleton augmentations. We use 2D skeleton heatmaps, which are fed to a video transformer for learning useful spatiotemporal and contextual features. Our method obtains higher accuracy and better robustness against missing and noisy keypoints, while showing superior performance in various fine-grained human activity understanding tasks. (c) We further fuse 2D skeleton heatmaps with RGB videos, establishing the state-of-the-art across all metrics and datasets.
@@
We are interested in self-supervised approaches for temporal video alignment, since acquiring fine-grained correspondence labels is generally hard and costly. Early methods [61, 23, 30, 49] attempt to match video frames directly. They optimize temporal alignment metrics such as dynamic time warping [18] and optimal transport [17] for self-supervised learning. Although the learned representations show promising generalization to downstream tasks, the results are far from being practical. Recently, CASA [41] proposes a context-aware self-supervised method for temporal video alignment, which operates on 3D skeletons (see Fig. 1(a)). It conducts 3D skeleton-based augmentations for self-supervised learning, while employing a transformer [70] to learn contextual features, yielding great results. Since CASA [41] directly processes 3D skeleton coordinates, its performance is greatly affected by human pose estimation errors such as noisy and missing keypoints.
We present a self-supervised temporal video alignment method, which relies on 2D skeletons (see Fig. 1(b)). Motivated by the great performance of [22] in skeleton-based action recognition, we convert sequences of 2D skeletons into sequences of heatmaps, where each joint is modeled as a Gaussian distribution, making it less sensitive to human pose estimation errors. Unlike CASA [41] which performs self-attention in the temporal domain only, we pass the heatmap sequences to our alignment network, which adopts a video transformer [2] to perform self-attention both in the spatial and temporal domains, yielding effective spatiotemporal and contextual features. Moreover, we propose simple various heatmap augmentation strategies based on 2D skeletons for self-supervised learning. As compared to CASA [41], our method based on 2D skeleton heatmaps obtains both higher accuracy and better robustness against missing and noisy keypoints despite lacking 3D information. Moreover, we show superior performance of our approach over previous methods in fine-grained human activity understanding tasks on three public datasets. Finally, utilizing both 2D skeleton heatmaps and RGB videos as inputs achieves the state-of-the-art across all metrics and datasets.
@@
  • We develop a multi-modality variant, which achieves the state-of-the-art on all metrics and datasets. To the best of our knowledge, our work is the first to exploit 2D skeleton heatmap inputs and the first to perform multi-modality fusion for temporal video alignment.
@@ 2 Related Work
Self-Supervised Learning. Considerable efforts have been made in developing pretext tasks with pseudo labels for training image-based self-supervised models. Examples include image colorization [42, 43], object counting [56, 50], predicting rotations [26], solving puzzles [9, 37], image inpainting [34], and image clustering [10, 11]. Recently, several works [12, 13, 28, 14] focus on designing data augmentation and contrastive learning strategies, which are important for learning effective representations. Unlike the above image-based methods, video-based methods, which exploit both spatial cues and temporal cues in videos, have attracted a great amount of interests recently. They learn representations by using pretext tasks such as future frame prediction [66, 71, 1, 21] and frame clustering [39], or leveraging temporal information such as temporal coherence [55, 82, 27], temporal order [53, 44, 25, 74], arrow of time [58, 73], and pace [5, 72, 77], or utilizing contrastive learning techniques [24, 33, 60, 20].
@@
Skeleton-Based Action Recognition. There exist notable interests in skeleton-based action recognition due to its compactness. A popular group of methods [76, 64, 7, 15, 29, 65] model the skeleton sequence as a spatiotemporal graph and rely on Graph Convolutional Networks (GCNs). Despite great results, their drawbacks lie in robustness to noisy/incomplete skeletons [81], scalability with multi-person actions, and fusion with other modalities [19]. Another group of methods convert the skeleton sequence into a 2D input (e.g., heatmap aggregation [16, 75, 3] and coordinate aggregation [36, 45, 51, 6]) or a 3D input (e.g., stacked pseudo images [32, 47] and cuboid [48]) and employ Convolutional Neural Networks (CNNs). These methods often suffer from information loss during aggregation. To address that, Duan et al. [22] convert each skeleton into a heatmap and pass the heatmap sequence, which preserves all information, to a CNN, and achieve superior results while alleviating the above issues of GCN-based methods. Inspired by their method for coarse-grained action recognition, we propose a self-supervised temporal video alignment framework for fine-grained tasks, e.g., action phase classification. In addition to video-level cues, fine-grained tasks further require frame-level cues and architectures/losses for extracting/utilizing them, which makes them more challenging.
@@ 3 Our Approach
Below we describe our main contribution, a video-based self-supervised learning method using 2D skeleton-based temporal video alignment as a pretext task. An overview of our approach is illustrated in Fig. 2.
Notations. Let us denote $\bm{S}=\{\bm{s}_{1},\bm{s}_{2},\dots,\bm{s}_{M}\}$ and $\bm{S}^{\prime}=\{\bm{s}^{\prime}_{1},\bm{s}^{\prime}_{2},\dots,\bm{s}^{\prime}_{N}\}$ as the original 2D skeleton sequence and its augmented sequence, with length $M$ and $N$ respectively. Each 2D skeleton has $K$ 2D joints, i.e., $\bm{s}_{i}\in\mathbb{R}^{K\times 2}$ and $\bm{s}^{\prime}_{j}\in\mathbb{R}^{K\times 2}$. The heatmaps obtained from $\bm{S}$ and $\bm{S}^{\prime}$ are written as $\bm{H}=\{\bm{h}_{1},\bm{h}_{2},\dots,\bm{h}_{M}\}$ and $\bm{H}^{\prime}=\{\bm{h}^{\prime}_{1},\bm{h}^{\prime}_{2},\dots,\bm{h}^{\prime}_{N}\}$, with $\bm{h}_{i}$ and $\bm{h}^{\prime}_{j}$ computed from $\bm{s}_{i}$ and $\bm{s}^{\prime}_{j}$ respectively. Next, let us represent the embedding function (up to the cross attention module) as $f_{\bm{\theta}}$, with learnable parameters $\bm{\theta}$. The embedding features of $\bm{H}$ and $\bm{H}^{\prime}$ are expressed as $\bm{U}=\{\bm{u}_{1},\bm{u}_{2},\dots,\bm{u}_{M}\}$ and $\bm{U}^{\prime}=\{\bm{u}^{\prime}_{1},\bm{u}^{\prime}_{2},\dots,\bm{u}^{\prime}_{N}\}$, with $\bm{u}_{i}=f_{\bm{\theta}}(\bm{h}_{i})$ and $\bm{u}^{\prime}_{j}=f_{\bm{\theta}}(\bm{h}^{\prime}_{j})$ respectively. Lastly, we denote the projection head function as $g_{\bm{\phi}}$, with learnable parameters $\bm{\phi}$. The latent features before the loss function are written as $\bm{Z}=\{\bm{z}_{1},\bm{z}_{2},\dots,\bm{z}_{M}\}$ and $\bm{Z}^{\prime}=\{\bm{z}^{\prime}_{1},\bm{z}^{\prime}_{2},\dots,\bm{z}^{\prime}_{N}\}$, with $\bm{z}_{i}=g_{\bm{\phi}}(\bm{u}_{i})$ and $\bm{z}^{\prime}_{j}=g_{\bm{\phi}}(\bm{u}^{\prime}_{j})$ respectively.
@@
In the following, we discuss the estimation of 2D skeletons $\bm{S}$ and their corresponding heatmaps $\bm{H}$ from a video sequence. An off-the-shelf 2D human pose estimator such as OpenPose [8] can be used to extract good quality 2D skeletons. For memory efficiency, a 2D skeleton is often stored as a set of 2D joint triplets $\{(x_{k},y_{k},c_{k})\}$, where $(x_{k},y_{k})$ are the 2D coordinates and $c_{k}$ is the (maximum) confidence score of the $k$-th joint. In practice, for fair comparison purposes, we use 2D skeletons that are projected (via orthographic projection) from 3D skeletons provided by CASA [41] and set $c_{k}=1.0$. We use orthographic projection since camera intrinsics/extrinsics are not available for all datasets (e.g., Penn Action). Our results with OpenPose 2D poses can be found in the supplementary material.
@@
Unlike CASA [41] which operates on 3D skeleton coordinates, our method relies on 2D skeleton heatmaps. Advantages of 2D skeleton heatmaps over 3D skeleton coordinates include: 1) Accurate 2D skeletons are generally easier to obtain, as accurate depths are not required (e.g., see Fig. 1 of [22]), 2) 2D skeleton heatmaps are more robust, as joints/limbs are modeled by Gaussian maps, 3) 2D skeleton heatmaps are applicable to advanced CNNs, whereas MLPs/GCNs are often used for 3D skeleton coordinates, and 4) It is convenient to fuse 2D skeleton heatmaps with other grid-like data. The above advantages contribute to our superior accuracy and robustness, as shown later in Sec. 4.
@@
Temporal Augmentation. We adjust the action speed by randomly selecting $N$ 2D skeletons out of the $M$ 2D skeletons in $\bm{S}$, resulting in $\bm{S}^{\prime}$.
Position Augmentation. We perturb the 2D joints in $\bm{S}$ by adding a random noise (sampled from a Gaussian distribution $\mathcal{N}(\nu)$ with a standard deviation $\nu$), yielding $\bm{S}^{\prime}$.
Orientation Augmentation. We pick a limb in $\bm{S}$ (e.g., arm, leg) and rotate it by a random angle (sampled from a Gaussian distribution $\mathcal{N}(\rho)$ with a standard deviation $\rho$), resulting in $\bm{S}^{\prime}$.
Horizontal Flipping. We flip the 2D joints in $\bm{S}$ from left to right and vice versa, yielding $\bm{S}^{\prime}$.
@@ 3.4 Multi-Modality Fusion
Representing skeletons as heatmaps makes it convenient to fuse with other modalities with grid structures such as RGB videos, optical flows, and depth maps. In this section, we extend our approach by using both skeleton heatmaps and RGB videos as inputs and performing late fusion, as demonstrated in Fig. 5. Our RGB+Pose model begins with separate encoders which respectively map the original skeleton heatmaps $\bm{H}$ and the augmented skeleton heatmaps $\bm{H}^{\prime}$ to features $\bm{U_{H}}$ and $\bm{U_{H}}^{\prime}$, and respectively map the original video frames $\bm{F}$ and the augmented video frames $\bm{F}^{\prime}$ to features $\bm{U_{F}}$ and $\bm{U_{F}}^{\prime}$. For heatmap augmentation, we apply the techniques in Sec. 3.2, while for video frame augmentation, we employ temporal augmentation, brightness augmentation, contrast augmentation, and horizontal flipping. Next, $\bm{U_{H}}$ and $\bm{U_{F}}$ are fused together to yield the combined features $\bm{U}$ for the original sequences, and $\bm{U_{H}}^{\prime}$ and $\bm{U_{F}}^{\prime}$ are fused together to yield the combined features $\bm{U}^{\prime}$ for the augmented sequences. Our fusion module includes concatenating the feature vectors and a linear layer to reduce the concatenated feature vector to the original size of 192. The fused features $\bm{U}$ and $\bm{U}^{\prime}$ are then fed to the projection heads to obtain the final features $\bm{Z}$ and $\bm{Z}^{\prime}$, which are used to perform matching and compute the loss as in Fig. 2. To train our RGB+Pose model, we perform two-stage training, i.e., we first train the encoders for each modality separately using the single-modality alignment loss, and then train them jointly with late fusion.
@@ 4 Experiments
Datasets. We evaluate on IKEA ASM [4], H2O [40], and Penn Action [79]. Following [30, 49, 41], we experiment on the Kallax Drawer Shelf videos of IKEA ASM [4], which demonstrates furniture assembly activities. H2O [40] is a hand-object interaction dataset. We evaluate on the Pouring Milk activity of H2O, following [41]. Penn Action [79] contains sporting activities. We use the same 13 actions of Penn Action as [23, 30, 49, 41]. These datasets are captured with diverse camera viewpoints, actors, backgrounds, and actions, as illustrated in Fig. 6. We use full-body poses for Penn Action [79] and IKEA ASM [4], hand poses for H2O [40], and the same training/testing splits as [23, 30, 49, 41].
@@
Evaluation Metrics. We compute four metrics on the validation set. We train the network on the training set and freeze it, and then train an SVM classifier or a linear regressor on top of the frozen features. We use the below metrics:
  • Phase Classification Accuracy is the percentage of correctly classified frames for fine-grained action recognition, achieved by training an SVM classifier on extracted features to predict phase labels.
  • Phase Progression measures the ability of the learned representations to predict the progress of an action over time, computed as the average R-squared measure between the ground truth progress values and those predicted by a linear regressor trained on the embeddings.
  • Kendall’s Tau is used to assess temporal alignment. It is in [-1,1], where 1 denotes perfect alignment and -1 denotes alignment in reverse order. It does not require any labels and assumes a strictly monotonic order of actions.
  • Average Precision is uses to assess fine-grained frame retrieval, computed as the proportion of retrieved frames with the same phase labels as the query frame. It is in [0,1], where 1 denotes perfect retrieval and vice versa.

### v6

# %TODO␣FINAL:␣Replace␣with␣your␣institution␣list.www.retrocausal.ai

@@ 1 Introduction
00footnotetext: indicates joint first author. {huy,ahmed,murad,hassan,andrey,zeeshan}@retrocausal.ai.
Figure 1: (a) The state-of-the-art method in self-supervised learning, i.e., CASA, uses 3D skeleton-based temporal video alignment as a pretext task and performs 3D skeleton augmentations. (b) Our approach relies on 2D skeleton-based temporal video alignment and conducts 2D skeleton augmentations. We use 2D skeleton heatmaps, which are fed to a video transformer for learning useful spatiotemporal and contextual features. Our method obtains higher accuracy and better robustness against missing and noisy keypoints, while showing superior performance in various fine-grained human activity understanding tasks. (c) We further fuse 2D skeleton heatmaps with RGB videos, establishing the state-of-the-art across all metrics and datasets.
@@
We are interested in self-supervised approaches for temporal video alignment, since acquiring fine-grained correspondence labels is generally hard and costly [23, 38]. Early methods [62, 23, 30, 50] attempt to match video frames directly. They optimize different temporal alignment metrics such as dynamic time warping [18] and optimal transport [17] for self-supervised learning. Although the learned representations show promising generalization to downstream tasks, the results are far from being practical. Recently, CASA [43] proposes a context-aware self-supervised method for temporal video alignment, which operates on 3D skeletons (see Fig. 1(a)). It conducts 3D skeleton-based augmentations for self-supervised learning, while employing a transformer [72] to learn contextual features, yielding great results. Since CASA [43] directly processes 3D skeleton coordinates, its performance is greatly affected by human pose estimation errors such as noisy and missing keypoints.
We present a self-supervised temporal video alignment method, which relies on 2D skeletons (see Fig. 1(b)). Motivated by the recent successes of [22, 34], we convert sequences of 2D skeletons into sequences of heatmaps, where each joint is modeled as a Gaussian distribution, making it less sensitive to human pose estimation errors. Unlike CASA [43] which performs self-attention in the temporal domain only, we pass the heatmap sequences to our alignment network, which adopts a video transformer [2] to perform self-attention both in the spatial and temporal domains, yielding effective spatiotemporal and contextual features. Moreover, we propose simple various heatmap augmentation strategies based on 2D skeletons for self-supervised learning. As compared to CASA [43], our method based on 2D skeleton heatmaps obtains both higher accuracy and better robustness against missing and noisy keypoints despite lacking 3D information. Moreover, we show superior performance of our approach over previous methods in fine-grained human activity understanding tasks on three public datasets. Finally, utilizing both 2D skeleton heatmaps and RGB videos as inputs achieves the state-of-the-art across all metrics and datasets.
@@
  • We develop a multi-modality variant, achieving the state-of-the-art on all metrics and datasets. To our best knowledge, our work is the first to exploit 2D skeleton heatmap inputs and perform multi-modality fusion for temporal video alignment.
@@ 2 Related Work
Self-Supervised Learning. Considerable efforts have been made in developing pretext tasks with pseudo labels for training image-based self-supervised models. Examples include image colorization [44, 45], object counting [57, 51], predicting rotations [26], solving puzzles [9, 39], image inpainting [35], and image clustering [10, 11]. Recently, several works [12, 13, 28, 14] focus on designing data augmentation and contrastive learning strategies, which are important for learning effective representations. Unlike the above image-based methods, video-based methods, which exploit both spatial cues and temporal cues in videos, have attracted a great amount of interests recently. They learn representations by using pretext tasks such as future frame prediction [67, 73, 1, 21] and frame clustering [41, 71], or leveraging temporal information such as temporal coherence [56, 84, 27], temporal order [54, 46, 25, 76], arrow of time [59, 75], and pace [5, 74, 79], or utilizing contrastive learning techniques [24, 33, 61, 20].
@@
Skeleton-Based Action Recognition. There exist notable interests in skeleton-based action recognition due to its compactness. A popular group of methods [78, 65, 7, 15, 29, 66] model the skeleton sequence as a spatiotemporal graph and rely on Graph Convolutional Networks (GCNs). Despite great results, their drawbacks lie in robustness to noisy/incomplete skeletons [83], scalability with multi-person actions, and fusion with other modalities [19]. Another group of methods convert the skeleton sequence into a 2D input (e.g., heatmap/coordinate aggregation [37, 16, 47, 52, 77, 6, 3]) or a 3D input (e.g., stacked pseudo images [32, 49]) and employ Convolutional Neural Networks (CNNs). These methods often suffer from information loss during aggregation. To address that, Duan et al. [22] convert each skeleton into a heatmap and pass the heatmap sequence, which preserves all information, to a CNN, and achieve superior results while alleviating the above issues of GCN-based methods. Inspired by their method for coarse-grained action recognition, we propose a self-supervised temporal video alignment framework for fine-grained tasks, e.g., action phase classification. In addition to video-level cues, fine-grained tasks further require frame-level cues and architectures/losses for extracting/utilizing them, which makes them more challenging.
@@ 3 Our Approach
Below we describe our main contribution, a self-supervised learning method using 2D skeleton-based temporal video alignment as a pretext task. Fig. 2 illustrates an overview.
Notations. Let us denote $\bm{S}=\{\bm{s}_{1},\bm{s}_{2},\dots,\bm{s}_{M}\}$ and $\bm{S}^{\prime}=\{\bm{s}^{\prime}_{1},\bm{s}^{\prime}_{2},\dots,\bm{s}^{\prime}_{N}\}$ as the original 2D skeleton sequence and its augmented sequence, with length $M$ and $N$ respectively. Each 2D skeleton has $K$ 2D joints, i.e., $\bm{s}_{i}\in\mathbb{R}^{K\times 2}$ and $\bm{s}^{\prime}_{j}\in\mathbb{R}^{K\times 2}$. The heatmaps obtained from $\bm{S}$ and $\bm{S}^{\prime}$ are written as $\bm{H}=\{\bm{h}_{1},\bm{h}_{2},\dots,\bm{h}_{M}\}$ and $\bm{H}^{\prime}=\{\bm{h}^{\prime}_{1},\bm{h}^{\prime}_{2},\dots,\bm{h}^{\prime}_{N}\}$, with $\bm{h}_{i}$ and $\bm{h}^{\prime}_{j}$ computed from $\bm{s}_{i}$ and $\bm{s}^{\prime}_{j}$ respectively. Next, let us represent the embedding function (up to the cross attention module) as $f_{\bm{\theta}}$, with learnable parameters $\bm{\theta}$. The embedding features of $\bm{H}$ and $\bm{H}^{\prime}$ are expressed as $\bm{U}=\{\bm{u}_{1},\bm{u}_{2},\dots,\bm{u}_{M}\}$ and $\bm{U}^{\prime}=\{\bm{u}^{\prime}_{1},\bm{u}^{\prime}_{2},\dots,\bm{u}^{\prime}_{N}\}$, with $\bm{u}_{i}=f_{\bm{\theta}}(\bm{h}_{i})$ and $\bm{u}^{\prime}_{j}=f_{\bm{\theta}}(\bm{h}^{\prime}_{j})$ respectively. Lastly, we denote the projection head function as $g_{\bm{\phi}}$, with learnable parameters $\bm{\phi}$, and the latent features before the loss function as $\bm{Z}=\{\bm{z}_{1},\bm{z}_{2},\dots,\bm{z}_{M}\}$ and $\bm{Z}^{\prime}=\{\bm{z}^{\prime}_{1},\bm{z}^{\prime}_{2},\dots,\bm{z}^{\prime}_{N}\}$, with $\bm{z}_{i}=g_{\bm{\phi}}(\bm{u}_{i})$ and $\bm{z}^{\prime}_{j}=g_{\bm{\phi}}(\bm{u}^{\prime}_{j})$ respectively.
@@
We discuss the estimation of 2D skeletons $\bm{S}$ and their corresponding heatmaps $\bm{H}$ from a video sequence. An off-the-shelf 2D human pose estimator such as OpenPose [8] can be used to extract good quality 2D skeletons. For memory efficiency, a 2D skeleton is often stored as a set of 2D joint triplets $\{(x_{k},y_{k},c_{k})\}$, where $(x_{k},y_{k})$ are the 2D coordinates and $c_{k}$ is the (maximum) confidence score of the $k$-th joint. In practice, for fair comparisons, we use 2D skeletons that are projected (via orthographic projection) from 3D skeletons provided by CASA [43] and set $c_{k}=1.0$. We use orthographic projection since camera intrinsics/extrinsics are not available for all datasets (e.g., Penn Action). Our results with OpenPose 2D poses can be found in the supplementary material.
@@
2D Skeleton Heatmaps vs. 3D Skeleton Coordinates. Unlike CASA [43] which operates on 3D skeleton coordinates, our method relies on 2D skeleton heatmaps. Advantages of 2D skeleton heatmaps over 3D skeleton coordinates include: 1) Accurate 2D skeletons are generally easier to obtain, as accurate depths are not required (e.g., see Fig. 1 of [22]), 2) 2D skeleton heatmaps are more robust, as joints/limbs are modeled by Gaussian maps, 3) 2D skeleton heatmaps are applicable to advanced CNNs, whereas MLPs/GCNs are often used for 3D skeleton coordinates, and 4) It is convenient to fuse 2D skeleton heatmaps with other grid-like data. The above advantages contribute to our superior accuracy and robustness, as shown later in Sec. 4.
@@
Temporal Augmentation. We adjust action speed by randomly selecting $N$ 2D skeletons out of $M$ 2D skeletons in $\bm{S}$, resulting in $\bm{S}^{\prime}$.
Position Augmentation. We perturb 2D joints in $\bm{S}$ by adding a random noise (sampled from a Gaussian distribution $\mathcal{N}(\nu)$ with a standard deviation $\nu$), yielding $\bm{S}^{\prime}$.
Orientation Augmentation. We pick a limb in $\bm{S}$ (arm or leg) and rotate it by a random angle (sampled from a Gaussian distribution $\mathcal{N}(\rho)$ with a standard deviation $\rho$), resulting in $\bm{S}^{\prime}$.
Horizontal Flipping. We flip 2D joints in $\bm{S}$ (left to right or vice versa), yielding $\bm{S}^{\prime}$.
@@ 3.4 Multi-Modality Fusion
Representing skeletons as heatmaps makes it convenient to fuse with other modalities with grid structures such as RGB videos, optical flows, and depth maps. In this section, we extend our approach by using both skeleton heatmaps and RGB videos as inputs and performing late fusion. Due to space constraints, an illustration of our RGB+Pose model architecture is included in the supplementary material. Our RGB+Pose model begins with separate encoders which respectively map the original skeleton heatmaps $\bm{H}$ and the augmented skeleton heatmaps $\bm{H}^{\prime}$ to features $\bm{U_{H}}$ and $\bm{U_{H}}^{\prime}$, and respectively map the original video frames $\bm{F}$ and the augmented video frames $\bm{F}^{\prime}$ to features $\bm{U_{F}}$ and $\bm{U_{F}}^{\prime}$. For heatmap augmentation, we apply the techniques in Sec. 3.2, while for video frame augmentation, we employ temporal augmentation, brightness augmentation, contrast augmentation, and horizontal flipping. Next, $\bm{U_{H}}$ and $\bm{U_{F}}$ are fused together to yield the combined features $\bm{U}$ for the original sequences, and $\bm{U_{H}}^{\prime}$ and $\bm{U_{F}}^{\prime}$ are fused together to yield the combined features $\bm{U}^{\prime}$ for the augmented sequences. Our fusion module includes concatenating the feature vectors and a linear layer to reduce the concatenated feature vector to the original size of 192. The fused features $\bm{U}$ and $\bm{U}^{\prime}$ are then fed to the projection heads to obtain the final features $\bm{Z}$ and $\bm{Z}^{\prime}$, which are used to perform matching and compute the loss as in Fig. 2. To train our RGB+Pose model, we perform two-stage training, i.e., we first train the encoders for each modality separately using the single-modality alignment loss, and then train them jointly with late fusion.
Differences with CASA [43]. There are several differences between our method and CASA [43], including: 1) Our method simply requires 2D skeletons (vs. 3D skeletons in CASA). 2) Our method performs self-attention both in spatial and temporal domains (vs. no spatial self-attention in CASA). 3) Our simple heatmap augmentations are based on 2D skeletons (vs. 3D skeletons in CASA). 4) Our method obtains both higher accuracy and better robustness than CASA. 5) Our multi-modality fusion further improves the performance (vs. no multi-modality fusion in CASA).
@@ 4 Experiments
Datasets. We evaluate on IKEA ASM [4], H2O [42], and Penn Action [81]. Following [30, 50, 43], we experiment on the Kallax Drawer Shelf videos of IKEA ASM [4], showing a furniture assembly activity. H2O [42] is a hand-object interaction dataset. We evaluate on the Pouring Milk activity of H2O, following [43]. Penn Action [81] contains sporting activities. We use the same 13 actions of Penn Action as [23, 30, 50, 43]. These datasets are captured with diverse camera viewpoints, actors, backgrounds, and actions, as illustrated in Fig. 5. We use full-body poses for Penn Action [81] and IKEA ASM [4], hand poses for H2O [42], and the same training/testing splits as [23, 30, 50, 43].
@@
Evaluation Metrics. We compute four metrics on the validation set, i.e., Phase Classification Accuracy, Phase Progression, Kendall’s Tau, and Average Precision. Please refer to [23, 30, 50, 43] for the definitions. We train the network on the training set and freeze it, and then train an SVM classifier or a linear regressor on top of the frozen features.